Show your work. 1) If the period of a pendulum is 12s, What is its frequency? (0.083Hz) What is its length? (35.8m)
Answer #1

I)Frequency=1/(time period)

Here ,time period=12 s.

So, frequency=1/12=0.083 Hz.

II) Time period=2π(l/g)^0.5, where l is length of the string and g is gravitational acceleration.

So,2π*(l/9.8)^0.5=12

=>l/9.8=[12/(2π)]^2=3.64756

=> l=3.64756*9.8=35.8 m.